i am doing quadratic equations how do i solve by factoring?
(x-1)^2 = 7

Answer by ankor@dixie-net.com(22740)

solve by factoring?
(x-1)^2 = 7
:
(x-1)(x-1) = 7
:
x^2 - 2x + 1 = 7
:
x^2 - 2x + 1 - 7 = 0
:
x^2 - 2x - 6 = 0
:
Problem here is, it will not factor.
:
solve it by completing the square:
(x - 1)^2 = 7
(x - 1) = +/-Sqrt(7)
x = 1 + Sqrt(7)
and
x = 1 - Sqrt(7)